If a parade 1 mi long is proceeding at 4 mph, how long will it take a runner jogging at 6 mph to travel from the front of the parade to the end of the parade?
Answer by checkley77(12844)

6-4=2 MPH. IS THE RELATIVE SPEED OF THE RUNNER.
D=RT
1=2/T
T=1/2
T=.5 HOURS TO PASS THE PARADE.